Please graph the solution set for the linear equation for x-3=y
Answer by ewatrrr(24785)


Hi


graphing:
y = x-3 |slope m = 1/1 and y-intercept is Pt(0,-3).
Plotting Pt(0,-3) and points (1,-2)& (3,0) determined using the slope m = 1/1
Note*** as y increases/decreases by 1, x increases/decreases by 1